Prove that

(v.del)v = 1/2 del v^2 - v×(del×v)

v is a vector
1
Expert's answer
2020-07-20T19:04:30-0400

SInce, (https://en.wikipedia.org/wiki/Del)

"{\displaystyle \nabla ({\vec {u}}\cdot {\vec {v}})=({\vec {u}}\cdot \nabla ){\vec {v}}+({\vec {v}}\cdot \nabla ){\vec {u}}+{\vec {u}}\times (\nabla \times {\vec {v}})+{\vec {v}}\times (\nabla \times {\vec {u}})}"

Now, put "\overrightarrow{u}=\overrightarrow{v}" ,we get

"{\displaystyle \nabla ({\vec {v}}\cdot {\vec {v}})=({\vec {v}}\cdot \nabla ){\vec {v}}+({\vec {v}}\cdot \nabla ){\vec {v}}+{\vec {v}}\times (\nabla \times {\vec {v}})+{\vec {v}}\times (\nabla \times {\vec {v}})}"

"\implies {\displaystyle \nabla ( {v^2})=2({\vec {v}}\cdot \nabla ){\vec {v}}+2{\vec {v}}\times (\nabla \times {\vec {v}})}\\
\implies{ ({\vec {v}}\cdot \nabla ){\vec {v}}= \frac{1}{2}{\displaystyle \nabla ( {v^2})}-{\vec {v}}\times (\nabla \times {\vec {v}})}"